OpenStudy (anonymous):

Solve the system of equations. 6x + 2y = 10 and -6x - 2y = 12

OpenStudy (anonymous):

@primeralph do u know how to solve this?

OpenStudy (anonymous):

-6x - 2y = 12 -2y -12 = 6x Sub 6x in the first equation: -2y - 12 + 2y = 10 -12 = 10 (N.A) No solution

OpenStudy (primeralph):

No solution.

OpenStudy (primeralph):